In the short run, a competitive firm has a production function,

Q = f(L) = 2.6667L0.75. The output price is $4 per unit and the wage is $5 per hour. Find the shortrun labor demand curve of the firm.


1
Expert's answer
2021-10-13T20:43:05-0400

"Q= f(L)= 2.6667L^{0.75}"

Wage,w =5

Output price, p =4.

Labor demand curve, "L= f(w,p)"


"MP_L= \frac{\delta Q}{\delta L}=\alpha AL^{\alpha -1}"

"= 0.75\times2.6667\times L^{0.75-1}"

"=2L^{0.75-1}"

"MP_L= \frac{W}{P}"

"2L^{-0.25}=\frac{5}{4}"

"L=5."

The short run labor demand will be 5.